Using Green's theorem,evaluate ∮c(x+y)dx+ x²dy, C is the triangle

with vertices at (0,0). (2,0) and (2, 4), taken in that order.


1
Expert's answer
2021-08-28T06:16:03-0400

The components of the vector field are

"P\n(\nx\n,\ny\n)\n=\nx^2\n,\nQ\n(\nx\n,\ny\n)\n=\nx\n+\ny"

Using the Green’s formula

"∬_
R
 
(\frac{
∂
Q}{
∂
x}
−
\frac{∂
P}{
∂
y}
)
d
x
d
y
=
∮_
C
 
P
d
x
+
Q
d
y"

we transform the line integral into the double integral:

"I
=
∮_
C
 
x^2
d
y
+
(
x
+
y
)
d
x
=
∬_
R
 
(
\frac{∂
(
x
+
y
)}{
∂
x}
−
\frac{∂
(
x
y
)}{
∂
y}
)
d
x
d
y
=
∬_
R
 
(
2x
−
1
)
d
x
d
y"

"\int _0^2\int _0^{2x}\left(2x-1\right)dydx\\
=\int _0^22x\left(2x-1\right)dx\\
=\frac{20}{3}"